How many grams of NaCN would you need to dissolve in enough water to make exactly 250 mL of solution with a pH of 10.00?

(Ka, HCN = 4.9 X 10^-10)


1
Expert's answer
2020-12-06T10:28:34-0500

pH + pOH = 14

pOH = 14 – 10 = 4

[OH-] = "10^{-4} \;M"

NaCN + H2O <=> HCN + OH-

To calculate the initial molarity of NaCN:

Let the concentration of NaCN be x such that

Equilibrium:

"[NaCN] = (x – 10^{-4}) \; M \\

[HCN] = 10^{-4} \;M \\

[OH^-] = 10^{-4} \;M \\

K_b = \frac{[HCN][OH^-]}{[NaCN]} \\

K_b = \frac{K_w}{K_a} \\

K_b = \frac{10^{-14}}{4.9 \times 10^{-10}} = 2 \times 10^{-5} \\

2 \times 10^{-5} = \frac{10^{-4} \times 10^{-4}}{(x – 10^{-4})} \\

x = 6 \times 10^{-4} \\

[NaCN] = 6 \times 10^{-4} \;M \\

n(NaCN) = 6 \times 10^{-4} \;M \times 0.250 \;L = 1.5 \times 10^{-4} \;mol \\

M(NaCN) = 49.01 \;g/mol \\

m(NaCN) = 49.01 \times 1.5 \times 10^{-4} = 7.35 \times 10^{-3} \;g"

Answer: "7.35 \times 10^{-3} \;g"